frank is older than beth. the sum of their age is 23 and the difference is 7. find frank and beths age
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

No similar or duplicate problems.

Answer by stanbon(75887)

frank is older than beth. the sum of their age is 23 and the difference is 7. find frank and beths age
-------
f + b = 23
f - b = 7
-----
2f = 30
frank = 15 yrs
-----
b = 23-f = 23-15 = 8 yrs
-------------------
Cheers,
Stan H.
------------